The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 46 in., find the width of the rectangle.
Answer by vertciel(183)

Hello there,
Perimeter of square (a rectangle is a square BTW) = sum of all sides
Let w be width.
So width = w and length = 2 + w since length is 2 inches longer than w.
Perimeter of square = w + (2 + w) + w + (2 + w) <--- There are 2 lengths and 2 widths in a square.
4w + 4 = 46
Find w and that's your width. Length is 2 + w so just add 2 to width and you have your answer in inches.